OpenStudy (anonymous):

Help here please, Triangles Can you show me how to find the missing sides and angles for ABC with a = 4.3, b=5.2, and c=7.5. I dont want the answer to these, I want to know how to solve it. Thank you:)

OpenStudy (inkyvoyd):

You arelooking for the law of cosines.

OpenStudy (inkyvoyd):

By the way, the pythagorean theorem is actually a special case of the law of cosines.

OpenStudy (inkyvoyd):

c^2=a^2+b^2-2ab cos C

OpenStudy (anonymous):

ok. onse sec, il solve it now

OpenStudy (anonymous):

c^2 = a^2 + b^2 - 2ab\cos C\,

OpenStudy (inkyvoyd):

a, b, and c are all interchangeable, for example a=b', b=c', c=a'

OpenStudy (anonymous):

a,b,c are all sides while A,B,C are angles

OpenStudy (anonymous):

one question. Why is it c² and not a². or it doesnt matter?

OpenStudy (inkyvoyd):

Ok.

OpenStudy (inkyvoyd):

For any arbitray triangle

OpenStudy (inkyvoyd):